OpenStudy (anonymous):

write Cos^2(x/2) - Sin(2x) in terms of Sin(x) and Cos(X) . ((hint use a double angle and half angle identity ))

OpenStudy (marco26):

 Then add them.

OpenStudy (anonymous):

i don't know how to do that can u show me?

OpenStudy (marco26):



OpenStudy (anonymous):

dat's the answer

OpenStudy (marco26):

\[\frac{ 1+cosx }{ 2 }-2sinxcosx=\frac{ 1+cosx-4sinxcosx}{2}=\frac{ 1+cosx(1-4sinx)}{2}\]